a particle is located in a two dimensional square potential well with absolutely, impenetrable walls (0‹x‹l, 0‹y‹l). find the probability of finding the particle within a region 0‹x‹l/3, 0‹y‹l/3 with lowest energy.


1
Expert's answer
2021-07-05T08:42:12-0400

"P=\frac S{S_0}=\frac{\frac l3\cdot \frac l3}{l\cdot l}=\frac 19."